OpenStudy (anonymous):

Solve n3 + 2n2 - 15n = 0 by factoring. Show the factored form of the equation, and the resulting solutions.

OpenStudy (anonymous):

n(n+5)(n-3)=0 it means the answers 0.-5,3

hartnn (hartnn):

n3+5n2-3n2-15n=0 n2(n+5)-3n(n+5)=0 (n2-3n)(n+5)=0 n(n-3)(n+5)=0 n=0,3,-5